OpenStudy (anonymous):

Can someone check my answer please? rad(x+1)+4=2x I got x=-5

OpenStudy (anonymous):

\[\sqrt{x+1}+4=2x\]

OpenStudy (anonymous):

\[2(\sqrt{x=1})=(2x-4)2\] \[x+1=4x+16\] \[-3x=15\] \[x=-5\]

OpenStudy (jdoe0001):

\(\bf (2x-4)^2\implies (2x-4)(2x-4) \ne 2^2x-4^2\)

OpenStudy (anonymous):

Okay so my answer is wrong? @jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Then what do I do?

OpenStudy (jdoe0001):

is a quadratic...so it should have 2 values

OpenStudy (anonymous):

How do I get the two values? I'm confused on where to start now

OpenStudy (jdoe0001):

\(\bf \sqrt{x+1}+4=2x\implies x+1=(2x-4)^2 \\ \quad \\ x+1=(2x)^2-2(2x)(4)+(4)^2\implies x+1=4x^2-16x+16 \\ \quad \\ 0=4x^2-17x+15\)

OpenStudy (nikato):

You messed up on your first/second step when you were trying to get rid of the radical sign You multiplied both sides by 2 When you should have squared both sides. Multiplying by 2 doesn't get rip do the radical sign

OpenStudy (jdoe0001):

in order to get rid of the root, you raise the term by THAT MUCH in this case, since the root is 2, you raise it by 2

OpenStudy (anonymous):

okay thanks that makes more sense! @jdoe0001

OpenStudy (jdoe0001):

yw